# Decision Quotient: A Regime-Sensitive Complexity Theory of Exact Relevance Certification

### 16 Mar 2026 v1

@@ 11 Conclusion
Within the encoding model of Section 2.4, we establish:
  • Static structure and regime matrix: in the static regime, the exact information content of the decision is the relevant-coordinate set itself: sufficient sets are exactly its supersets, so exact minimization reduces to relevance containment and is coNP-complete rather than a genuine $\Sigma_{2}^{\textsf{P}}$ search problem, while Anchor-Sufficiency remains $\Sigma_{2}^{\textsf{P}}$-complete. In the stochastic regime, preservation (stochastic sufficiency) has a polynomial-time base query in the explicit-state model with proved bridges to static sufficiency and the optimizer quotient; the minimum and anchor preservation variants are likewise polynomial-time in explicit state, inherit coNP-completeness and $\Sigma_{2}^{\textsf{P}}$-completeness under full support, and admit support-sensitive obstruction and bridge lemmas beyond full support. The obstruction to general-distribution classification is structural: preservation quantifies over all states, while stochastic analysis is driven by probability-weighted fibers. Zero-probability states can witness violations that the distribution itself never observes, so full-support results do not transfer directly. The support-sensitive obstruction lemmas in the artifact isolate the exact combinatorial mechanism. Decisiveness (stochastic decisiveness) has a polynomial-time base query in the explicit-state model and is PP-hard under succinct encoding; the anchor and minimum decisiveness queries are PP-hard and lie in $\textsf{NP}^{\textsf{PP}}$, with those oracle-class memberships proved in the paper text and their finite witness/checking cores mechanized. In the sequential regime, sufficiency, minimum, and anchor queries are PSPACE-complete.
  • Encoding-sensitive contrast for exact certification: logarithmic-size sufficient sets admit polynomial-time algorithms in the explicit regime, while linear-size sufficient sets inherit ETH-conditioned exponential lower bounds in the succinct regime.
  • Structural and downstream consequences: the paper derives witness and approximation obstructions, exact-simplification hardness transfers, and a direct reliability corollary for exact certifiers in the hard regime: no exact certifier can simultaneously be sound, complete, and polynomial-budgeted on that regime. A simplicity-tax framing shows that omitted exact relevance is externalized rather than eliminated.
  • Structured tractability: tractable subcases across two tiers: simple edge cases (single action, bounded state space, strict dominance, constant optimal set) admitting immediate checks, and structurally nontrivial cases (bounded actions, separable utility, low tensor rank, tree structure, bounded treewidth, coordinate symmetry) each removing a distinct hardness source with a mechanized reduction or decision procedure.
Taken together, these results show that exact relevance certification is a distinct complexity-theoretic object in its own right: computing $\text{Opt}(s)$ on one state can be easy even when proving coordinate irrelevance is hard, and the difficulty of that certification problem depends sharply on whether one is ruling out counterexamples, preserving optimizers under coarser conditioning, checking decisiveness of conditional fibers, or reasoning about temporally structured optimizer behavior.
@@
The implications are direct hardness transfers. Configuration simplification is an instance of sufficiency checking, so exact behavior-preserving minimization has no general-purpose polynomial-time worst-case solution unless $\textsf{P}=\textsf{coNP}$. The same issue appears in real pipeline orchestration settings, where a nested configuration object controls well selection, output layout, and downstream materialization. There too the practical question is whether one can hide part of the configuration interface without changing the induced compilation or execution decisions. Our results show that this is an exact certification problem whose worst-case difficulty depends on whether one is in the static, stochastic, or sequential regime. Reliable exact certification is likewise regime-limited: under $\textsf{P}\neq\textsf{coNP}$, the trilemma of exact certification rules out any solver that is simultaneously sound, complete on the full hard regime, and polynomial-budgeted. And once the canonical relevant-coordinate set $R(\mathcal{D})$ is fixed, any compressed interface merely partitions that relevance into centrally handled and externalized parts. Compression therefore relocates exact relevance rather than erasing it; under the per-site cost model, the unresolved part grows linearly with deployment scale.
@@
The proof artifact mechanically checks the main reduction constructions, hardness packages, finite deciders, search procedures, tractability statements, step-counting wrappers, the stochastic sufficiency bridge package, the explicit-state decision procedures for the preservation family, the full-support inheritance results, the new support-sensitive obstruction and bridge lemmas for preservation, the decisiveness upper bounds, and the finite witness/checking schemata for the stochastic existential classification (L: DC91-99, OU8-12, EH2-10.) Full oracle-machine formalization in Lean remains outside the current state of formalized complexity theory; the paper proves the oracle-class memberships in the text following standard conventions, while the artifact provides independent verification of the finite combinatorial core. This places the mechanization in the emerging category of problem-specific certified reduction infrastructure.
@@
Two immediate directions remain. First, the reduction infrastructure can be integrated more tightly with general-purpose formalized complexity libraries. Second, the artifact boundary can be made even cleaner by continuing to package the existing finite-decision results into more uniform summary interfaces without changing the paper’s complexity claims. On the complexity side, the remaining gap is now sharply localized: stochastic decisiveness already has succinct hardness and a verified witness/checking core for the oracle-class memberships proved in the paper text, but a fully standard oracle-machine formalization of the resulting oracle-class placement and of the $\textsf{NP}^{\textsf{PP}}$ boundary for the anchor and minimum decisiveness queries remains outside the current repository; on the preservation side, Conjecture 4.1 and Open Problem 4.2 sharpen the remaining open terrain to the full general-distribution and succinct-encoding complexity of the minimum and anchor preservation variants. The mechanized artifact already exposes the finite combinatorial core and support-sensitive lemmas that isolate this obstruction; see lean_handle_ids_auto.tex and claim_mapping_auto.tex for handles such as OU12, DC96, and DC98, which are appropriate citations when framing the open problem.
@@ Acknowledgments
Generative AI tools (including Codex, Claude Code, Augment, Kilo, and OpenCode) were used throughout this manuscript, across all sections (Abstract, Introduction, theoretical development, proof sketches, applications, conclusion, and appendix) and across all stages from initial drafting to final revision. The tools were used for boilerplate generation, prose and notation refinement, /structure cleanup, translation of informal proof ideas into candidate formal artifacts (Lean/), and repeated adversarial reviewer-style critique passes to identify blind spots and clarity gaps.
@@
Appendix A Lean Handle Index
The table below lists the auto-generated Lean handle identifiers cited inline in the paper.
ID Handle ID Handle
AB2 DecisionProblem.surjective_abstraction_factors_or_erases DecisionQuotient/AbstractionCollapse.lean CC1 ClaimClosure.DQ8 DecisionQuotient/ClaimClosure.lean
CR1 DecisionQuotient.ConfigReduction.config_sufficiency_iff_behavior_preserving DecisionQuotient/Hardness/ConfigReduction.lean CT12 DecisionQuotient.Physics.ClaimTransport.physical_bridge_bundle DecisionQuotient/Physics/ClaimTransport.lean
DC19 StochasticSequential.stochastic_anchor_sufficient_of_stochastic_sufficient DecisionQuotient/StochasticSequential/Quotient.lean DC23 StochasticSequential.sequential_anchor_sufficient_of_sequential_sufficient DecisionQuotient/StochasticSequential/Basic.lean
DC28 StochasticSequential.reduceTQBF_correct_anchor DecisionQuotient/StochasticSequential/PolynomialReduction.lean DC29 StochasticSequential.reduceTQBF_to_sequential_anchor_reduction DecisionQuotient/StochasticSequential/PolynomialReduction.lean
DC40 StochasticSequential.reduceMAJSATPureAnchor_correct DecisionQuotient/StochasticSequential/PolynomialReduction.lean DC41 StochasticSequential.reduceMAJSAT_to_pure_stochastic_anchor_reduction DecisionQuotient/StochasticSequential/PolynomialReduction.lean
DC42 StochasticSequential.stochastic_anchor_check_pp_hard DecisionQuotient/StochasticSequential/PolynomialReduction.lean DC44 StochasticSequential.sequential_anchor_check_pspace_hard DecisionQuotient/StochasticSequential/PolynomialReduction.lean
DC45 StochasticSequential.stochastic_sufficiency_pp_hard DecisionQuotient/StochasticSequential/PolynomialReduction.lean DC46 StochasticSequential.sequential_sufficiency_pspace_hard DecisionQuotient/StochasticSequential/PolynomialReduction.lean
DC47 StochasticSequential.stochastic_minimum_sufficiency_pp_hard DecisionQuotient/StochasticSequential/PolynomialReduction.lean DC48 StochasticSequential.sequential_minimum_sufficiency_pspace_hard DecisionQuotient/StochasticSequential/PolynomialReduction.lean
DC49 StochasticSequential.sequentialMinimalSufficient_iff_relevant DecisionQuotient/StochasticSequential/Basic.lean DC50 StochasticSequential.sequentialRelevantSet_is_minimal DecisionQuotient/StochasticSequential/Basic.lean
DC57 StochasticSequential.fiberDecisionProblem_sufficient DecisionQuotient/StochasticSequential/Basic.lean DC62 StochasticSequential.countedStochasticSufficiencySearch_spec DecisionQuotient/StochasticSequential/Computation.lean
DC63 StochasticSequential.countedStochasticSufficiencySearch_steps DecisionQuotient/StochasticSequential/Computation.lean DC64 StochasticSequential.countedStochasticAnchorSearch_spec DecisionQuotient/StochasticSequential/Computation.lean
DC65 StochasticSequential.countedStochasticAnchorSearch_steps DecisionQuotient/StochasticSequential/Computation.lean DC66 StochasticSequential.countedSequentialSufficiencySearch_spec DecisionQuotient/StochasticSequential/Computation.lean
DC67 StochasticSequential.countedSequentialSufficiencySearch_steps DecisionQuotient/StochasticSequential/Computation.lean DC68 StochasticSequential.countedSequentialAnchorSearch_spec DecisionQuotient/StochasticSequential/Computation.lean
DC69 StochasticSequential.countedSequentialAnchorSearch_steps DecisionQuotient/StochasticSequential/Computation.lean DC70 DecisionQuotient.static_sufficiency_inP_explicit DecisionQuotient/ExplicitStateMembership.lean
DC71 DecisionQuotient.static_anchor_inP_explicit DecisionQuotient/ExplicitStateMembership.lean DC72 StochasticSequential.stochastic_sufficiency_inP_explicit DecisionQuotient/StochasticSequential/Computation.lean
DC73 StochasticSequential.stochastic_anchor_inP_explicit DecisionQuotient/StochasticSequential/Computation.lean DC74 StochasticSequential.sequential_sufficiency_inP_explicit DecisionQuotient/StochasticSequential/Computation.lean
DC75 StochasticSequential.sequential_anchor_inP_explicit DecisionQuotient/StochasticSequential/Computation.lean DC76 DecisionQuotient.explicit_state_inP_summary DecisionQuotient/ExplicitStateMembership.lean
DC80 DecisionQuotient.staticSufficiency_counted_search_witness DecisionQuotient/AlgorithmComplexity.lean DC81 DecisionQuotient.static_query_search_matrix DecisionQuotient/AlgorithmComplexity.lean
DC82 DecisionQuotient.finite_search_summary DecisionQuotient/ComplexityMain.lean DC91 StochasticSequential.countedStochasticPreservationSearch_spec DecisionQuotient/StochasticSequential/Computation.lean
DC92 StochasticSequential.countedStochasticPreservationSearch_steps DecisionQuotient/StochasticSequential/Computation.lean DC93 StochasticSequential.stochastic_preservation_inP_explicit DecisionQuotient/StochasticSequential/Computation.lean
DC94 StochasticSequential.stochastic_preservation_implies_static_sufficiency DecisionQuotient/StochasticSequential/Basic.lean DC95 StochasticSequential.static_sufficiency_implies_stochastic_preservation_of_full_support DecisionQuotient/StochasticSequential/Basic.lean
DC96 StochasticSequential.static_sufficiency_iff_stochastic_preservation_of_full_support DecisionQuotient/StochasticSequential/Basic.lean DC97 StochasticSequential.stochasticDecisionEquiv_iff_decisionEquiv_of_preservation DecisionQuotient/StochasticSequential/Quotient.lean
DC98 StochasticSequential.stochasticDecisionEquiv_iff_decisionEquiv_of_full_support DecisionQuotient/StochasticSequential/Quotient.lean DC99 StochasticSequential.stochasticEquivSetoid_eq_decisionSetoid_of_full_support DecisionQuotient/StochasticSequential/Quotient.lean
DC100 StochasticSequential.benchmark_escalation_summary DecisionQuotient/StochasticSequential/Hierarchy.lean DN1 DecisionProblem.stateDecisionNoise_iff_same_quotient DecisionQuotient/DecisionNoise.lean
DN2 DecisionProblem.decisionNoise_iff_not_relevant DecisionQuotient/DecisionNoise.lean DN5 DecisionProblem.decisionNoise_iff_condIndep DecisionQuotient/DecisionNoise.lean
DP1 DecisionQuotient.DecisionProblem.minimalSufficient_iff_relevant DP2 DecisionQuotient.DecisionProblem.relevantSet_is_minimal
DP6 ClaimClosure.DP6 DecisionQuotient/ClaimClosure.lean DP7 ClaimClosure.DP7 DecisionQuotient/ClaimClosure.lean
DQ1 ClaimClosure.DQ1 DecisionQuotient/ClaimClosure.lean DQ2 ClaimClosure.DQ2 DecisionQuotient/ClaimClosure.lean
DQ3 ClaimClosure.DQ3 DecisionQuotient/ClaimClosure.lean DQ6 ClaimClosure.DQ6 DecisionQuotient/ClaimClosure.lean
DQ7 ClaimClosure.DQ7 DecisionQuotient/ClaimClosure.lean DQ9 DecisionQuotient.HardnessDistribution.simplicityTax_conservation DecisionQuotient/HardnessDistribution.lean
DQ10 DecisionQuotient.IntegrityCompetence.alwaysAbstain_integrity DecisionQuotient/IntegrityCompetence.lean DQ11 DecisionQuotient.SetCoverInstance.min_sufficient_iff_set_cover DecisionQuotient/Hardness/SetCoverReduction.lean
DQ12 DecisionQuotient.StochasticSequential.bounded_horizon_tractable DecisionQuotient/StochasticSequential/Tractability.lean DQ13 DecisionQuotient.StochasticSequential.bounded_support_tractable DecisionQuotient/StochasticSequential/Tractability.lean
DQ14 DecisionQuotient.StochasticSequential.fully_observable_tractable DecisionQuotient/StochasticSequential/Tractability.lean DQ15 DecisionQuotient.StochasticSequential.product_distribution_tractable DecisionQuotient/StochasticSequential/Tractability.lean
DQ16 DecisionQuotient.StochasticSequential.product_enables_transfer DecisionQuotient/StochasticSequential/CrossRegime.lean DQ18 DecisionQuotient.StochasticSequential.static_simpler_than_stochastic DecisionQuotient/StochasticSequential/CrossRegime.lean
DQ19 DecisionQuotient.StochasticSequential.stochastic_simpler_than_sequential DecisionQuotient/StochasticSequential/CrossRegime.lean DQ20 DecisionQuotient.all_coordinates_necessary_of_not_tautology DecisionQuotient/Hardness/MinSufficientApproximation.lean
DQ25 DecisionQuotient.counted_min_sufficient_inapproximability_conditional DecisionQuotient/Hardness/ApproximationHardness.lean DQ26 DecisionQuotient.dichotomy_conditional DecisionQuotient/ClaimClosure.lean
DQ27 DecisionQuotient.eth_lower_bound_informal DecisionQuotient/Hardness/ETH.lean DQ28 DecisionQuotient.exact_certainty_inflation_under_hardness_core DecisionQuotient/ClaimClosure.lean
DQ30 DecisionQuotient.hard_family_all_coords_core DecisionQuotient/ClaimClosure.lean DQ31 DecisionQuotient.integrity_resource_bound_for_sufficiency DecisionQuotient/ClaimClosure.lean
DQ32 DecisionQuotient.min_sufficient_factor_inapprox_of_set_cover_factor_inapprox DecisionQuotient/Hardness/ApproximationHardness.lean DQ34 DecisionQuotient.min_sufficient_set_cover_equiv DecisionQuotient/Hardness/ApproximationHardness.lean
DQ36 DecisionQuotient.minsuff_conp_complete_conditional DecisionQuotient/ClaimClosure.lean DQ38 DecisionQuotient.no_exact_claim_admissible_under_hardness_core DecisionQuotient/ClaimClosure.lean
DQ39 DecisionQuotient.singleton_gate_sufficient_of_tautology DecisionQuotient/Hardness/MinSufficientApproximation.lean DQ40 DecisionQuotient.sufficiency_conp_complete_conditional DecisionQuotient/ClaimClosure.lean
DQ42 DecisionQuotient.sufficient_means_factorizable DecisionQuotient/Information.lean DQ43 DecisionQuotient.tautology_decidable_from_factor_approx DecisionQuotient/Hardness/MinSufficientApproximation.lean
DQ44 DecisionQuotient.tractable_subcases_conditional DecisionQuotient/ClaimClosure.lean DQ45 DecisionQuotient.tractable_tree_core DecisionQuotient/ClaimClosure.lean
DQ51 DecisionQuotient.bounded_actions_tractable DecisionQuotient/Summary.lean DQ56 DecisionQuotient.orbitType_count_bound DecisionQuotient/Tractability/Dimensional.lean
DQ57 DecisionQuotient.separable_utility_tractable DecisionQuotient/Summary.lean DQ58 DecisionQuotient.stochastic_objective_bridge_can_fail_on_sufficiency DecisionQuotient/ClaimClosure.lean
DQ62 DecisionQuotient.transition_coupled_bridge_can_fail_on_sufficiency DecisionQuotient/ClaimClosure.lean DQ63 DecisionQuotient.tree_structure_tractable DecisionQuotient/Summary.lean
DQ72 DecisionQuotient.bounded_actions_complexity DecisionQuotient/Tractability/BoundedActions.lean DQ74 DecisionQuotient.sufficiency_poly_separable DecisionQuotient/Tractability/SeparableUtility.lean
DQ75 DecisionQuotient.tensor_contraction_tractable DecisionQuotient/Tractability/SeparableUtility.lean DQ76 DecisionQuotient.low_rank_utility_admits_factored_computation DecisionQuotient/Tractability/SeparableUtility.lean
DQ77 DecisionQuotient.low_rank_tractability DecisionQuotient/Tractability/SeparableUtility.lean DQ78 DecisionQuotient.sufficiency_poly_tree_structured DecisionQuotient/Tractability/TreeStructure.lean
DQ79 DecisionQuotient.csp_treewidth_tractable DecisionQuotient/Tractability/TreeStructure.lean DQ80 DecisionQuotient.sufficiency_reduces_to_interaction_csp DecisionQuotient/Tractability/TreeStructure.lean
DQ81 DecisionQuotient.bounded_treewidth_tractability DecisionQuotient/Tractability/TreeStructure.lean DQ82 DecisionQuotient.orbitType_eq_iff DecisionQuotient/Tractability/Dimensional.lean
DQ83 DecisionQuotient.symmetric_optimalActions_orbit_invariant DecisionQuotient/Tractability/Dimensional.lean DQ84 DecisionQuotient.sufficiency_reduces_to_cross_orbit_check DecisionQuotient/Tractability/Dimensional.lean
DQ85 DecisionQuotient.symmetric_sufficiency_complexity_bound DecisionQuotient/Tractability/Dimensional.lean EH1 StochasticSequential.existential_anchor_source_fits_np_over_ppstyle_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean
EH2 StochasticSequential.existential_anchor_hard_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean EH3 StochasticSequential.existential_anchor_query_family_hard_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean
EH4 StochasticSequential.existential_anchor_np_over_ppstyle_hard_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean EH5 StochasticSequential.existential_anchor_query_family_np_over_ppstyle_hard_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean
EH6 StochasticSequential.existential_anchor_query_family_np_over_ppstyle_complete_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean EH7 StochasticSequential.existential_decisiveness_complement_np_over_ppstyle_hard_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean
EH8 StochasticSequential.existential_decisiveness_complement_np_over_ppstyle_complete_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean EH9 StochasticSequential.existential_decisiveness_query_family_np_over_ppstyle_hard_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean
EH10 StochasticSequential.existential_decisiveness_query_family_np_over_ppstyle_complete_honest DecisionQuotient/StochasticSequential/ExistentialHardness.lean EX1 Examples.pomdp_reduction_to_preservation DecisionQuotient/Examples/PreservationExamples.lean
EX2 Examples.hyperparam_reduction_to_static EX3 Examples.toy_full_opt_s1_contains_a DecisionQuotient/Examples/ConcreteExamples.lean
HD4 DecisionQuotient.HardnessDistribution.complete_model_dominates_after_threshold DecisionQuotient/HardnessDistribution.lean HD5 DecisionQuotient.HardnessDistribution.gap_conservation_card DecisionQuotient/HardnessDistribution.lean
HD14 DecisionQuotient.HardnessDistribution.linear_lt_exponential_plus_constant_eventually DecisionQuotient/HardnessDistribution.lean HD23 DecisionQuotient.HardnessDistribution.hardness_is_irreducible_required_work DecisionQuotient/HardnessDistribution.lean
HD26 DecisionQuotient.HardnessDistribution.totalExternalWork_eq_n_mul_gapCard DecisionQuotient/HardnessDistribution.lean IC30 DecisionQuotient.IntegrityCompetence.exactCertaintyInflation_iff_no_exact_competence DecisionQuotient/IntegrityCompetence.lean
IC32 DecisionQuotient.IntegrityCompetence.integrity_forces_abstention DecisionQuotient/IntegrityCompetence.lean IC33 DecisionQuotient.IntegrityCompetence.integrity_not_competent_of_nonempty_scope DecisionQuotient/IntegrityCompetence.lean
IC34 DecisionQuotient.IntegrityCompetence.integrity_resource_bound DecisionQuotient/IntegrityCompetence.lean OU1 StochasticSequential.stochastic_anchor_query_fits_np_over_ppstyle DecisionQuotient/StochasticSequential/OracleUpperBounds.lean
OU2 StochasticSequential.stochastic_minimum_query_fits_np_over_ppstyle DecisionQuotient/StochasticSequential/OracleUpperBounds.lean OU6 StochasticSequential.stochastic_anchor_inP_explicit_via_witness_schema DecisionQuotient/StochasticSequential/OracleUpperBounds.lean
OU7 StochasticSequential.stochastic_minimum_inP_explicit_via_witness_schema DecisionQuotient/StochasticSequential/OracleUpperBounds.lean OU8 StochasticSequential.fitsCoNPOverPPStyle_of_no_witness DecisionQuotient/StochasticSequential/OracleUpperBounds.lean
OU9 StochasticSequential.stochastic_decisiveness_query_fits_conp_over_ppstyle DecisionQuotient/StochasticSequential/OracleUpperBounds.lean OU10 StochasticSequential.stochastic_decisiveness_complement_fits_np_over_ppstyle DecisionQuotient/StochasticSequential/OracleUpperBounds.lean
OU11 StochasticSequential.stochastic_decisiveness_scoped_oracle_bounds DecisionQuotient/StochasticSequential/OracleUpperBounds.lean OU12 StochasticSequential.stochastic_preservation_explicit_summary DecisionQuotient/StochasticSequential/OracleUpperBounds.lean
PA3 Physics.AnchorChecks.stochastic_anchor_check_iff_exists_anchor_singleton DecisionQuotient/Physics/AnchorChecks.lean QT1 DecisionProblem.quotient_is_coarsest DecisionQuotient/Quotient.lean
QT7 DecisionProblem.quotient_has_unique_factorization DecisionQuotient/Quotient.lean WD1 DecisionQuotient.checking_witnessing_duality_budget DecisionQuotient/WitnessCheckingDuality.lean
WD2 DecisionQuotient.no_sound_checker_below_witness_budget DecisionQuotient/WitnessCheckingDuality.lean WD3 DecisionQuotient.checking_time_ge_witness_budget DecisionQuotient/WitnessCheckingDuality.lean

### 31 Mar 2026 v2

@@ XII Conclusion
The core classification is Theorem VI.1, with cross-encoding separation in Theorem VII.1, reliability limits in Theorem IX.6, and structured tractability in Theorem VIII.1. Together, these results show that exact relevance certification is computationally distinct from pointwise optimizer evaluation and depends sharply on regime (Theorem VI.1). The optimizer quotient remains the governing structure (Theorem II.15), and the remaining conjecture and open problem mark the boundary left unresolved.
@@
The implications are direct hardness transfers. Configuration simplification is an instance of sufficiency checking, so exact behavior-preserving minimization has no general-purpose polynomial-time worst-case solution unless $\mathrm{P}=\mathrm{coNP}$. The same issue appears in real pipeline orchestration settings, where a nested configuration object controls well selection, output layout, and downstream materialization. There too the practical question is whether one can hide part of the configuration interface without changing the induced compilation or execution decisions. Our results show that this is an exact certification problem whose worst-case difficulty depends on whether one is in the static, stochastic, or sequential regime. Reliable exact certification is likewise regime-limited: under $\mathrm{P}\neq\mathrm{coNP}$, the trilemma of exact certification rules out any solver that is simultaneously sound, complete on the full hard regime, and polynomial-budgeted. And once the relevant-coordinate set $R(\mathcal{D})$ is fixed, any compressed interface merely partitions that relevance into centrally handled and externalized parts. Compression therefore relocates exact relevance rather than erasing it; under the per-site cost model, the unresolved part grows linearly with deployment scale.
@@
The proof artifact mechanically checks the main reduction constructions, hardness packages, finite deciders, search procedures, tractability statements, step-counting wrappers, the stochastic sufficiency bridge package, the explicit-state decision procedures for the preservation family, the full-support inheritance results, the new support-sensitive obstruction and bridge lemmas for preservation, the decisiveness upper bounds, and the finite witness/checking schemata for the stochastic existential classification 848484: DC91-99, OU8-12, EH2-10. Full oracle-machine formalization in Lean remains outside the current state of formalized complexity theory; the paper proves the oracle-class memberships in the text following standard conventions, while the artifact provides independent verification of the finite combinatorial core. This places the mechanization in the emerging category of problem-specific certified reduction infrastructure.
@@
Two immediate directions remain. First, the reduction infrastructure can be integrated more tightly with general-purpose formalized complexity libraries. Second, the artifact boundary can be made even cleaner by continuing to package the existing finite-decision results into more uniform summary interfaces without changing the paper’s complexity claims. On the complexity side, the remaining gap is now sharply localized: stochastic decisiveness already has succinct hardness and a verified witness/checking core for the oracle-class memberships proved in the paper text, but a fully standard oracle-machine formalization of the resulting oracle-class placement and of the $\textsf{NP}^{\textsf{PP}}$ boundary for the anchor and minimum decisiveness queries remains outside the current repository; on the preservation side, Conjecture IV-D and Open Problem IV-D sharpen the remaining open terrain to the full general-distribution and succinct-encoding complexity of the minimum and anchor preservation variants. The mechanized artifact already exposes the finite combinatorial core and support-sensitive lemmas that isolate this obstruction; see the supplementary material for representative Lean handles when framing the open problem. 858585: OU12, DC96, DC98
@@ Acknowledgments
Generative AI tools (including Codex, Claude Code, Augment, Kilo, and OpenCode) were used throughout this manuscript, across all sections (Abstract, Introduction, theoretical development, proof sketches, applications, conclusion, and appendix) and across all stages from initial drafting to final revision. The tools were used for boilerplate generation, prose and notation refinement, /structure cleanup, translation of informal proof ideas into candidate formal artifacts (Lean/), and repeated adversarial critique passes to identify blind spots and clarity gaps.
@@
Acknowledgments
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC).
Appendix A Artifact Index and Supplementary Tables
The full Lean-handle ledger and claim-to-handle mapping are provided in the supplementary PDF (and archived artifact at https://doi.org/10.5281/zenodo.19057595). We omit those long tables from the main manuscript body.